OpenStudy (anonymous):

Find dy/dx of x^2*y+x*y^2=8

OpenStudy (anonymous):

Use implicit differentiation: (Using multiplication rule for derivatives) 2xy + y'x^2 + y^2 + 2xyy' = 0

OpenStudy (anonymous):

Put y' on one side: y'(x^2 + 2xy) = -y^2 -2xy

OpenStudy (anonymous):

y' = (-y^2 -2xy) / (x^2 + 2xy)

OpenStudy (anonymous):

Thanks dude!

OpenStudy (anonymous):

No problem :)